OpenStudy (marco26):

Find the differential equation of all tangent lines of the parabola y=x^2

OpenStudy (marco26):

My guess: y=x^2 The derivative of this function is the slope of its tangent line, so: y'=2x --> slope of the tangent line y=mx+b Since y =x^2, and m=y' x^2=y'x+b Differentiating wrt x will give: 2x=y'+xy'' ---> answer Is that right?